# Differentially Private Clustering in Data Streams

### v2

@@ Appendix A Proofs for Bicriteria Approximation
We set the failure probability of DP-HH (see Theorem 6) to be $\xi:=\frac{1}{k^{2}}$. For $\mathbf{c}\in G_{\ell}$, define ${\mathcal{E}}_{\mathbf{c}}:={\mathcal{E}}_{1}\wedge{\mathcal{E}}_{2}\wedge{\mathcal{E}}_{3}$ where ${\mathcal{E}}_{1}$ is the event that DP-HH algorithm is correct on all instances, ${\mathcal{E}}_{2}$ is the event that there are no collisions between $\mathbf{c}$ and other good cells, and ${\mathcal{E}}_{3}$ is the event that there exists a hash bucket that contains only $\mathbf{c}$, and if $N_{\ell,\mathbf{c}}:=y$, then the size of that hash bucket is $\leq 2(y+\frac{n_{G_{\ell}}}{40k})$. We know that $\Pr[{\mathcal{E}}_{1}]\geq 1-\xi=1-\frac{1}{k^{2}}$, by the accuracy guarantee of DP-HH algorithm, $\Pr[{\mathcal{E}}_{2}]\geq 1-\frac{1}{k^{2}}$ by Claim 2 (and boosting the success probability), and $\Pr[{\mathcal{E}}_{3}]\geq 1-\frac{1}{k^{2}}$, by Claim 3 (and boosting the success probability). Thus by a union bound, we have that for a fixed $\mathbf{c}\in G_{\ell}$, $\Pr[{\mathcal{E}}_{\mathbf{c}}]\geq 1-\frac{3}{k^{2}}$. By taking a union bound over all $4k$ good cells, with probability at least $1-\frac{12}{k}$, the claim holds. ∎
Proof of Claim 5.
@@
Recall that $N^{(t)}_{\ell,\mathbf{c}}$ is the number of points in cell $\mathbf{c}$ at time step $t$. Because $x_{t}$ is covered in $G_{\ell}$, this means that we only need to care about the points covered by a good cell $\mathbf{c}\in G_{\ell}$ but $\mathbf{c}\not\in H_{\ell,t}$. Using Claim 4, we know that if $N^{(t)}_{\ell,\mathbf{c}}\geq\frac{\theta n_{G_{\ell}}}{20k}$, and $N^{(t)}_{\ell,\mathbf{c}}\geq\frac{2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$, then with probability at least $1-\frac{12}{k}$, we have that $\mathbf{c}\in H_{\ell,t}$.
If $\mathbf{c}\in G_{\ell}$, and $\mathbf{c}\not\in H_{\ell,t}$, this means that either (1) $N^{(t)}_{\ell,\mathbf{c}}<\frac{2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$, or, (2) $\mathbf{c}$ is not an $\theta$-HH. For case (1), since there are $4k$ such good cells, the total number of uncovered points in such cells are $\frac{4k\cdot 2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$. For case (2), this means that $N^{(t)}_{\ell,\mathbf{c}}<\frac{\theta n_{G_{\ell}}}{20k}$. Again, since there are $4k$ such good cells, the total number of uncovered points in such cells are $<\frac{\theta n_{G_{\ell}}}{20k}\cdot 4k<\theta n_{G_{\ell}}$. ∎

### v3

@@ Appendix E Missing Proofs from Appendix B
We set the failure probability of DP-HH (see Theorem 10) to be $\xi:=\frac{1}{k^{2}}$. For $\mathbf{c}\in G_{\ell}$, define ${\mathcal{E}}_{\mathbf{c}}:={\mathcal{E}}_{1}\wedge{\mathcal{E}}_{2}\wedge{\mathcal{E}}_{3}$ where ${\mathcal{E}}_{1}$ is the event that DP-HH algorithm is correct on all instances, ${\mathcal{E}}_{2}$ is the event that there are no collisions between $\mathbf{c}$ and other good cells, and ${\mathcal{E}}_{3}$ is the event that there exists a hash bucket that contains only $\mathbf{c}$, and if $N_{\ell,\mathbf{c}}:=y$, then the size of that hash bucket is $\leq 2(y+\frac{n_{G_{\ell}}}{40k})$. We know that $\Pr[{\mathcal{E}}_{1}]\geq 1-\xi=1-\frac{1}{k^{2}}$, by the accuracy guarantee of DP-HH algorithm, $\Pr[{\mathcal{E}}_{2}]\geq 1-\frac{1}{k^{2}}$ by Proposition 2 (and boosting the success probability), and $\Pr[{\mathcal{E}}_{3}]\geq 1-\frac{1}{k^{2}}$, by Proposition 3 (and boosting the success probability). Thus by a union bound, we have that for a fixed $\mathbf{c}\in G_{\ell}$, $\Pr[{\mathcal{E}}_{\mathbf{c}}]\geq 1-\frac{3}{k^{2}}$. By taking a union bound over all $4k$ good cells, with probability at least $1-\frac{12}{k}$, the claim holds. ∎
Proof of Proposition 5.
@@
Recall that $N^{(t)}_{\ell,\mathbf{c}}$ is the number of points in cell $\mathbf{c}$ at time step $t$. Because $x_{t}$ is covered in $G_{\ell}$, this means that we only need to care about the points covered by a good cell $\mathbf{c}\in G_{\ell}$ but $\mathbf{c}\not\in H_{\ell,t}$. Using Proposition 4, we know that if $N^{(t)}_{\ell,\mathbf{c}}\geq\frac{\theta n_{G_{\ell}}}{20k}$, and $N^{(t)}_{\ell,\mathbf{c}}\geq\frac{2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$, then with probability at least $1-\frac{12}{k}$, we have that $\mathbf{c}\in H_{\ell,t}$.
If $\mathbf{c}\in G_{\ell}$, and $\mathbf{c}\not\in H_{\ell,t}$, this means that either (1) $N^{(t)}_{\ell,\mathbf{c}}<\frac{2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$, or, (2) $\mathbf{c}$ is not an $\theta$-HH. For case (1), since there are $4k$ such good cells, the total number of uncovered points in such cells are $\frac{4k\cdot 2\log(\Lambda)\log^{2}(k)}{\varepsilon\eta}\operatorname*{poly}(\log(\frac{T\cdot k\cdot 2^{\ell}}{\theta\eta}))$. For case (2), this means that $N^{(t)}_{\ell,\mathbf{c}}<\frac{\theta n_{G_{\ell}}}{20k}$. Again, since there are $4k$ such good cells, the total number of uncovered points in such cells are $\frac{\theta n_{G_{\ell}}}{20k}\cdot 4k<\theta n_{G_{\ell}}$. ∎